The Sisyphean Company has a bond outstanding with a face value of $5,000 that reaches maturity in 10 years. The bond certificate indicates that the stated coupon rate for this bond is 8.2% and that the coupon payments are to be made semiannually.

Assuming that this bond trades for $4,461, then the YTM for this bond is closest to:

a. 13.9%

b. 11.9%

c. 7.9%

d. 9.9%
Answer #1

Answer: Option d is correct

Face value=$5000
Coupon rate=8.2%
Semiannual coupon rate=8.2%/2=0.041
Semiannual coupon payment=(Semiannual coupon rate)*(Face value)=(0.041)*(5000)=205
Time period=10 years
As the coupon payments are made semiannually, the number of periosa=10*2=20

Now, we can determine the YTM using excel

205 1 Face value 5000 2 Payment 3 Number of periods 4 Present value -4461 5 YTM= 4.96% 6 Formula used: RATE(V3,12,14,v1)

So, the semiannual YTM=4.96%
Annual value of YTM =4.96%*2=9.92% or 9.9% (Rounded to one decimal place)

Note: As the present value is a cash outflow, we have taken it as negative in excel